OpenStudy (anonymous):

Find the limit of the function algebraically. limit as x approaches negative five of quantity x squared minus twenty five divided by quantity x plus five. Would the limit = -5?

OpenStudy (anonymous):

can you draw/type out the limit?

OpenStudy (anonymous):

Yes @peachpi

OpenStudy (anonymous):

\[\lim_{x \rightarrow -5}\frac{ x^2-25 }{ x+5}\]

OpenStudy (anonymous):

ok. can you factor the numerator?

OpenStudy (anonymous):

(x+5)(x-5)

OpenStudy (anonymous):

So the (x + 5) cancel and you have \[\lim_{x \rightarrow -5}x-5\] Now plug in -5 for x to get the limit

OpenStudy (anonymous):

Oh so -10?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Could you help me with one more?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Two triangles can be formed with the given information. Use the Law of Sines to solve the triangles. A = 56°, a = 16, b = 17 @peachpi

OpenStudy (anonymous):

\[\frac{ \sin A }{ a }=\frac{ \sin B }{ b }\] Solve for B to get the 2nd angle of the first triangle

OpenStudy (anonymous):

I did that and got b = .88 which isn't an answer choice.

OpenStudy (anonymous):

That's the answer choices I'm given.

OpenStudy (anonymous):

(sin 56°)/16 = (sin B)/17 sin B = (17 sin 56°)/16 sin B = 0.88 Oh, I see you. You need to take the inverse sine to find the angle. Use the sin^-1 button on your calculator \[B = \sin^{-1} 0.88 \] B = 61.74°

OpenStudy (anonymous):

Oh ok, so would the answer be B then?

OpenStudy (anonymous):

either B or D

OpenStudy (anonymous):

Oh...so how can I decide which?

OpenStudy (anonymous):

use law of sines again to test C. We know C = 62.3° in this triangle, so solve for c (sin 56°)/16 = (sin 62.3°)/c

OpenStudy (anonymous):

c = 17.1, so yeah B

OpenStudy (anonymous):

Ok thanks! Just did it on my calculator lol.

OpenStudy (anonymous):

you're welcome